OpenStudy (anonymous):

A radio station has a broadcast area in the shape of a circle with equation x2 + y2 = 5,625, where the constant represents square miles. a. Graph the equation and state the radius in miles. b. What is the area of the region in which the broadcast from the station can be picked up?

OpenStudy (anonymous):

put into y=mx+b form.

OpenStudy (anonymous):

How would I do that?

OpenStudy (anonymous):

well i am gonna guess that the equation is x^2+y^2=5625 so i would start by square rooting the whole equation.

OpenStudy (anonymous):

So x^2+y^2=31640625?

OpenStudy (anonymous):

no no no, square root..\[\sqrt{x^2+y^2}=\sqrt{5625}\]

OpenStudy (anonymous):

Oh. O.o Srry bout that then.

OpenStudy (anonymous):

no you are fine, so what would you get when you do the step above?

OpenStudy (anonymous):

x+y=75

OpenStudy (anonymous):

so then subtract x to get it in y=mx+b so you would have y=-x+75 go ahead and graph that

OpenStudy (anonymous):

WAIT! nevermind its a circle! crap crap crap nevermind! so sorry!

OpenStudy (anonymous):

lol s'okay

OpenStudy (anonymous):

the equation of a circle is x^2 + y^2 = r^2 compare this with your equation adn you will get the radius r

OpenStudy (anonymous):

*the general equation of a circle with radius 'r' and center in (a,b) is : (x-a)^2+(y-b)^2=r^2 compare this with your equation and find a,b and r , then you can graph it easily

OpenStudy (anonymous):

How would I substitute in what I have to this equation though?

OpenStudy (anonymous):

(x-a)^2 = x^2 then what is the value of a? and so on for b and r

OpenStudy (anonymous):

r^2=5625 --> then r = ?

OpenStudy (anonymous):

do you understand?